A scientist has two solutions, which she has labeled Solution A and Solution B. Each contains salt. She knows that Solution A is
70%
salt and Solution B is
95%
salt. She wants to obtain
30
ounces of a mixture that is
80%
salt. How many ounces of each solution should she use?

Answer by mananth(16946)

percent ---------------- quantity
salt solution I 70.00% ---------------- x ounces
salt solution II 95.00% ------ 30 - x ounces
Mixture 80.00% ---------------- 30
Total 30 ounces
70.00% x + 95.00% ( 30 - x ) = 80.00% * 30
70 x + 95 ( 30 - x ) = 2400
70 x + 2850 - 95 x = 2400
70 x - 95 x = 2400 - -2850
-25 x = -450
/ -25
x = 18 ounces 70.00% salt solution I
12 ounces 95.00% salt solution II

m.ananth@hotmail.ca